A space module is in an elliptical orbit around a planet.The maximum distance from the center of the planet to the module is 300 kilometers,and the minimum distance is 200 kilometers.If a model of this orbit is graphed with the center of the ellipse at the origin,write the equation that represents the path of the space module.
Answer by lwsshak3(11628)

A space module is in an elliptical orbit around a planet.The maximum distance from the center of the planet to the module is 300 kilometers,and the minimum distance is 200 kilometers.If a model of this orbit is graphed with the center of the ellipse at the origin,write the equation that represents the path of the space module.
..
Standard form for ellipse: x^2/a^2+y^2/b^2=1, a>b
For given problem:
length of major axis=600 km=2a
a=300
a^2=[(3)(10)^2]^2=(9)(10)^4
..
length of minor axis=400 km=2b
b=200
b^2=[(2)(10)^2]^2=(4)(10)^4
..
Equation: For path of module
x^2/(9)(10)^4+y2/(4)(10)^4=1